What is the ratio of the salt & acid concentration if the solution has a pH of 4.5 & pKa of 4.76?


1
Expert's answer
2022-02-01T18:25:01-0500

pH"=" pKa "+" log "\frac{[Salt]}{[Acid]}"


"4.5=4.76+log\>\frac{[Salt]}{[Acid]}"


Salt: Acid"=0.5495"